OpenStudy (ambermarie151):

Find the discriminant and the number of real roots for this equation. x2 + 2x + 8 = 0 A. 32; one real root B. –28; no real roots C. –28; one real root D. 32; two real roots

jhonyy9 (jhonyy9):

do you know the formula for discriminant ?

jhonyy9 (jhonyy9):

D = b^2 -4ac do you know the value of a,b and c ?

OpenStudy (ambermarie151):

Isn't like b2-4ac

jhonyy9 (jhonyy9):

yes this is

OpenStudy (ambermarie151):

A=1 b=2 c=8

jhonyy9 (jhonyy9):

yes this is correct

jhonyy9 (jhonyy9):

so just yous these value of a ,b and c and calculi the discriminant value

jhonyy9 (jhonyy9):

this is easy courage please

OpenStudy (ambermarie151):

2^2-4(1)(8) 4-32 =28?

jhonyy9 (jhonyy9):

4-32 = ? not 28

OpenStudy (ambermarie151):

It would be -28 ?

jhonyy9 (jhonyy9):

yes -28 so what is less than zero do you know the case when the discriminant is less than zero what roots has the quadratic ?

jhonyy9 (jhonyy9):

do you learned it ?

OpenStudy (ambermarie151):

It's would be one real ?

OpenStudy (ambermarie151):

Or because it's negative it would have no real

jhonyy9 (jhonyy9):

no there are 3 cases discriminant let D D > 0 there are two real ,different roots D = 0 there are two real equal roots D < 0 ?

OpenStudy (ambermarie151):

I'm so confused

jhonyy9 (jhonyy9):

yes your last answer is right sure

jhonyy9 (jhonyy9):

so when the discriminant is less than zero so the quadratic have not real roots ok. ?

OpenStudy (ambermarie151):

Okay

jhonyy9 (jhonyy9):

Or because it's negative it would have no real this your answer is right hope do you understand it now ok. ?

OpenStudy (ambermarie151):

Yes I do Thank you !

jhonyy9 (jhonyy9):

welcome was my pleasure and anytime good luck bye bye